# Jarne Verhaeghe IDLab

### v1

@@ Abstract
Understanding the dose-response relation between a continuous treatment and the outcome for an individual can greatly drive decision-making, particularly in areas like personalized drug dosing and personalized healthcare interventions. Point estimates are often insufficient in these high-risk environments, highlighting the need for uncertainty quantification to support informed decisions. Conformal prediction, a distribution-free and model-agnostic method for uncertainty quantification, has seen limited application in continuous treatments or dose-response models. To address this gap, we propose a novel methodology that frames the causal dose-response problem as a covariate shift, leveraging weighted conformal prediction. By incorporating propensity estimation, conformal predictive systems, and likelihood ratios, we present a practical solution for generating prediction intervals for dose-response models. Additionally, our method approximates local coverage for every treatment value by applying kernel functions as weights in weighted conformal prediction. Finally, we use a new synthetic benchmark dataset to demonstrate the significance of covariate shift assumptions in achieving robust prediction intervals for dose-response models.
@@ 1 Introduction
How can we determine the optimal dose for a patient to ensure the best therapeutic outcome? What is the impact of discounts in an online store on sales? What impact does $CO_{2}$ concentration have on local climates? At the core of each of these questions lies a shared causal idea: understanding the dose-response relation under continuous treatments to inform decision-making. In many cases, these decisions bear significant consequences, where relying solely on point estimates may be insufficient (Feuerriegel et al., 2024). Particularly in high-stakes situations, augmenting predictions with uncertainty quantification (UQ) can significantly improve decision-making processes (Feuerriegel et al., 2024). For instance, while the estimated causal effect of a continuous treatment may appear positive, prediction intervals could suggest a largely negative outcome for a specific individual. Such insights are crucial for deciding interventions. To tackle this, conformal prediction (CP) offers a robust solution for UQ, being both distribution-free and model-agnostic, with formal coverage guarantees (Vovk et al., 2022).
In this work, we seek to extend CP to UQ in dose-response models, aiming to aid decision-makers with more informed estimates to tackle such questions. We introduce a novel approach for deriving prediction intervals in the continuous treatment setting using weighted conformal prediction by combining propensity estimation with weighted conformal predictive systems. Furthermore, with the aid of a novel synthetic benchmark, we show how viewing the problem as a covariate shift approach provides coverage across all treatment values to help create more individualized dose-response curves.
@@ 2 Background
In this paper we expand upon the potential outcomes framework introduced in Rubin (2005), otherwise known as the Rubin framework to accommodate continuous treatments. Consider a continuous treatment variable $T\in[t_{L},t_{U}]$ with a lower bound $t_{L}$ and upper bound $t_{U}$, observed covariates $X$, and potential outcomes $Y(t)\in\mathbb{R}$ representing the outcome that would be observed under treatment level $t$. The Conditional Average Dose-Response Function (CADRF) is defined as $\nu(x,t)=E[Y(t)|X=x]$, the expected value over the Individual Dose-Response Functions (IDRF) for all individuals with observed $X$. Similar to Conditional Average Treatment Effects (CATE), to estimate the CADRF we make the following standard assumptions (Rubin, 2005; Hirano & Imbens, 2004):
  • Unconfoundedness: $Y(t)\perp\!\!\!\perp T|X,\,\forall t\in T$. This assumption states that, conditional on the observed covariates, the treatment assignment is independent of the potential outcomes. In other words, there are no unobserved confounders that influence both the treatment assignment and the outcome.
  • Overlap or positivity: $0<P(T=t|X=x)<1,\,\forall t\in T$ with $x\in X$. The overlap assumption ensures that for every covariate value $x$, there is a positive probability of receiving any treatment level. This is crucial for estimating treatment effects across the entire range of treatment levels.
  • Consistency: $Y=Y(t)$ with probability $1$. This assumption links the observed outcomes to the potential outcomes, stating that the observed outcome is equal to the potential outcome corresponding to the treatment received.
Quantifying the IDRF requires observing the $Y(t)$ for all possible treatment values. These treatment values are all counterfactuals and thus impossible to observe as we only can observe $Y$ for a single treatment value $t$ at a time. Furthermore for estimating the CADRF, likewise with CATE estimation, the distribution of the treatment assignment can bias the estimation (Hirano & Imbens, 2004). This distribution of the treatment assignment is called the propensity distribution, which was initially defined for binary treatments. Hirano & Imbens (2004) introduced the generalized propensity score (GPS) for continuous treatments that aims to unbias the CATE estimation for continuous treatments. The GPS is defined as $\pi(t_{i}|x)=f_{T|X}(T=t_{i}|X=x)$, which is the evaluation of $T=t_{i}$ on the conditional probability density function $T|X$ (Hirano & Imbens, 2004). If the treatment is independent of $X$, i.e. there are no confounders that influence treatment assignment, then $f_{T|X}$ is equal for all possible $X$. Furthermore, the treatment assignment is considered uniformly assigned between lower $t_{L}$ and upper $t_{U}$ possible treatment if $f_{T|X}$ represents the density function of the uniform distribution between $t_{L}$ and $t_{U}$. The GPS can then be used to mimic the randomly assigned treatment to estimate the unbiased CADRF (Wu et al., 2024).
@@
The estimated $\widehat{IDRF}$ can also be seen as follows: $\widehat{IDRF}=\nu(x,t)+\epsilon_{a,IDRF}(x,t)+\epsilon_{e,IDRF}(x,t)$. The aleatoric uncertainty is symbolized by $\epsilon_{a,IDRF}(x,t)$ created by the inherent variability between individuals having the same covariates. $\epsilon_{e,IDRF}(x,t)$ symbolises the epistemic uncertainty coming from model specification and finite samples. Estimating both uncertainties creates the opportunity to estimate the ranges of the $\widehat{IDRF}$:
Problem Definition
To accurately estimate the $\widehat{IDRF}$ for all possible treatment values we require correctly estimating both uncertainties for all treatment values equally, or more formally; for a specific significance level $\alpha$, lower treatment bound $t_{L}$, upper treatment bound $t_{U}$, and covariates $X$, we require prediction intervals $C(t,X)$ such that
@@ 3 Related Work
Propensity score methods, introduced by Rosenbaum & Rubin (1983), have become widespread in causal inference, especially in observational studies. These methods aim to balance confounders across treatment groups, reducing bias in treatment effect estimates. Hirano & Imbens (2004) generalized this propensity score to continuous instead of binary treatments, introducing the generalized propensity score and building the foundation for causal inference with continuous exposures. Wu et al. (2024) used the generalized propensity score for matching continuous treatments to debias the treatment assignment and more accurately estimate the average dose-response curve for all treatment values. Other approaches adapt machine learning techniques to dose-response modelling. For instance, Athey et al. (2019) developed generalized random forests for heterogeneous treatment effect estimation, adaptable to continuous treatments.
@@
To provide UQ, this work adapts conformal prediction. Conformal prediction is a model-agnostic method introduced by Vovk et al. (2022) that constructs prediction intervals with guaranteed finite-sample coverage under distribution-free assumptions. Conformal prediction uses conformity scores to assess uncertainty. Various improvements, such as the adaptive version by Romano et al. (2019), have increased the flexibility and applicability to even heteroscedastic settings. Additionally, Lei et al. (2018) and Papadopoulos et al. (2002) introduced split conformal prediction, significantly improving computational efficiency. For scenarios involving covariate or distribution shifts, Tibshirani et al. (2019) introduced weighted conformal prediction to ensure coverage under mismatched training and testing data distributions, with additional work by Gibbs & Candes (2021; 2024) and Barber et al. (2023). By reweighting the calibration samples similar to weighted conformal prediction,  Guan (2023) introduced localized conformal prediction where the prediction intervals are determined by calibration samples localized around the test sample. Vovk et al. (2019) also introduced conformal predictive systems (CPS); an extension of full conformal prediction that allows extracting predictive distributions instead of prediction intervals. More recently, Jonkers et al. (2024a) combined previous concepts, introducing weighted conformal predictive systems to also account for covariate shifts.
@@
In causal inference, conformal prediction has mainly been applied to binary treatments. For instance, Lei & Candès (2021) were among the first to apply conformal prediction to treatment effects estimation in randomized experiments and confounded or observational data. Jonkers et al. (2024b) and Alaa & Ahmad (2024) extended this approach to the potential outcomes framework, providing uncertainty to quantify individual treatment effects. However, the use of conformal prediction in continuous treatment settings remains largely unexplored. Schröder et al. (2024) proposed a conformal prediction framework for prediction intervals of treatment effects for continuous treatment interventions. However, their approach mainly covers single-treatment interventions and is computationally intensive, requiring optimization per confidence level, treatment, and sample where they provide prediction intervals for a single treatment value. For a more in-depth analysis of Schröder et al. (2024), see Appendix C.
@@ 4.1 Introduction to Conformal Prediction
Before delving into our proposed method, we provide a formal introduction to conformal prediction (Jonkers et al., 2024a; Tibshirani et al., 2019). Conformal prediction offers a powerful method for constructing prediction intervals with guaranteed finite-sample coverage under distribution-free assumptions (Vovk et al., 2022). The key insight of conformal prediction lies in its use of a nonconformity measure to quantify the degree to which a new observation differs from previously observed data.
@@
for a pre-specified significance level $\alpha\in(0,1)$ where the probability is calculated over the points $Z_{i},i=1,...,n$
@@
Finally, we construct the prediction interval containing all $y$ where (Jonkers et al., 2024a)
@@
Tibshirani et al. (2019) presented conformal prediction slightly differently by using quantile functions instead, which will be more convenient for weighted conformal prediction later on. Tibshirani et al. (2019) defines the $1-\alpha$ quantile function as follows, where $F_{R}(y)$ represents the distribution of nonconformity scores $R_{i}^{y}$ consisting of a sum of point masses $\delta_{a}$ with mass at $a$ where $R^{y}\sim F_{R}(y)$ (Tibshirani et al., 2019). $F_{R}(y)$ can then be used to calculate probabilities:
@@
This procedure guarantees that $P(y_{n+1}\in\hat{C}(X_{n+1}))\geq 1-\alpha$ for any exchangeable distribution of the data and any choice of nonconformity measure (Tibshirani et al., 2019).
@@ 4.1.1 Inductive Conformal Prediction
The previously mentioned conformal prediction approach is computationally heavy as it requires fitting $n\cdot\#\{\mathbb{R}\}+1$ estimators $\hat{\mu}$. Inductive or split conformal prediction (ICP), introduced by Papadopoulos et al. (2002), tackles this computation issue by splitting the training sequence $Z_{1:n}=\{Z_{1},...,Z_{n}\}$ into two sets: the proper training set $Z_{1:m}=\{Z_{1},...,Z_{m}\}$ and the calibration set $Z_{m+1:n}=\{Z_{m+1},...,Z_{n}\}$. A single regression model $\hat{\mu}$ is fit on the proper training set while the nonconformity scores (e.g., $R_{i}=|y_{i}-\hat{\mu}(X_{i})|,i=m+1,...,n$) are generated from the calibration set. These scores are sorted in descending order denoted as $R_{1}^{*},...,R_{n-m}^{*}$. Then, for a new sample with features $X_{n+1}$, a point prediction is made $\hat{y}_{n+1}=\hat{\mu}(X_{n+1})$. Finally, given a target coverage of $1-\alpha$, the prediction interval becomes
@@
where $s=\lfloor\alpha(n-m+1)\rfloor$ represents the $1-\alpha$ quantile of the ordered nonconformity set with size $n-m$ (Jonkers et al., 2024a).
@@ 4.1.2 Weighted Conformal Prediction
Evaluating and requiring coverage guarantees for the dose-response model at all possible treatment values changes the test distribution compared to the training distribution. In the training data, all treatment values are sampled according to their (conditional) training distribution, which can be determined by other variables in the case of confounding. However, every treatment value is possible in testing, and thus, every treatment sample can be sampled. This mimics sampling a new test sample with the treatment value from a uniform distribution, which can be vastly different from the treatment distribution in the training data. Standard conformal prediction only guarantees coverage if the joint distribution of the new sample $Z_{n+1}$ and $Z_{1:n}$ remains the same under permutations, which is called the exchangeability assumption (Vovk et al., 2022; Tibshirani et al., 2019). This issue is called covariate shift; The features $X_{n+1}$ come from a different distribution compared to $X_{1:n}$, while the relation between $X$ and $y$ remains the same. More formally: $X_{i}\sim P_{X},\ i=1,...,n$ and $X_{n+1}\sim\tilde{P}_{X}$ where $\tilde{P}_{X}\neq P_{X}$ while $y_{i}\sim P_{Y|X},\ i=1,...,n$.
Weighted conformal prediction provides a solution to tackle this issue (Tibshirani et al., 2019). However, their main assumption is that the likelihood ratio between the training $P_{X}$ and the test covariate distribution $\tilde{P}_{X}$ is known, defined as
@@
The rationale is that they reweight the distribution of nonconformity scores $F_{R}(y)$ to make the nonconformity scores more exchangeable with the test population by using the following weights in equation 7 (Tibshirani et al., 2019):
@@
Consequently, these weights adjust the distribution of nonconformity scores to give more weight to nonconformity scores that are more likely in the test set and vice versa while in standard conformal prediction, every $R_{i}$ has equal weight. Also, note that the weights $p^{w}(x)$ are normalized, cancelling out any constant terms resulting in $w(x)$ being proportional to $w(x)\propto\frac{d\tilde{P}(x)}{dP(x)}$. An extension to split weighted conformal prediction can be done similarly as in section 4.1.1 (Tibshirani et al., 2019).

### v2

# Conformal Prediction for Dose-Response Models with Continuous Treatments

@@ Abstract
Understanding the dose-response relation between a continuous treatment and the outcome for an individual can greatly drive decision-making, particularly in areas like personalized drug dosing and personalized healthcare interventions. Point estimates are often insufficient in these high-risk environments, highlighting the need for uncertainty quantification to support informed decisions. Conformal prediction, a distribution-free and model-agnostic method for uncertainty quantification, has seen limited application in continuous treatments or dose-response models. To address this gap, we propose a novel methodology that frames the causal dose-response problem as a covariate shift, leveraging weighted conformal prediction. By incorporating propensity estimation, conformal predictive systems, and likelihood ratios, we present a practical solution for generating prediction intervals for dose-response models. Additionally, our method approximates local coverage for every treatment value by applying kernel functions as weights in weighted conformal prediction. Finally, we use a new synthetic and semi-synthetic benchmark dataset to demonstrate the significance of covariate shift assumptions in achieving robust prediction intervals for counterfactual dose-response models.
@@ 1 Introduction
How can we determine the optimal dose for a patient to ensure the best therapeutic outcome? What is the impact of discounts in an online store on sales? What impact does $CO_{2}$ concentration have on local climates? At the core of each of these questions lies a shared causal idea: understanding the dose-response relation under continuous treatments to inform decision-making. In many cases, these decisions bear significant consequences, where relying solely on point estimates may be insufficient (feuerriegel_causal_2024). Particularly in high-stakes situations, augmenting predictions with uncertainty quantification (UQ) can significantly improve decision-making processes (feuerriegel_causal_2024). For instance, while the estimated causal effect of a continuous treatment may appear positive, prediction intervals could suggest a largely negative outcome for a specific individual. Such insights are crucial for deciding interventions. To tackle this, conformal prediction (CP) offers a robust solution for UQ, being both distribution-free and model-agnostic, with formal coverage guarantees (vovk_algorithmic_2022).
In this work, we seek to extend CP to UQ in dose-response models, aiming to aid decision-makers with more informed estimates to tackle such questions. We introduce a novel approach for deriving prediction intervals in the continuous treatment setting using weighted conformal prediction by combining propensity estimation with weighted conformal predictive systems. Furthermore, with the aid of a novel synthetic and semi-synthetic benchmark, we show how viewing the problem as a covariate shift approach provides coverage across all treatment values to help create more individualized dose-response curves.
@@ 2 Background
In this paper we expand upon the potential outcomes framework introduced in rubin_causal_2005, otherwise known as the Rubin framework to accommodate continuous treatments. Consider a continuous treatment variable $T\in[t_{L},t_{U}]$ with a lower bound $t_{L}$ and upper bound $t_{U}$, observed covariates $X$, and potential outcomes $Y(t)\in\mathbb{R}$ representing the outcome that would be observed under treatment level $t$. The Conditional Average Dose-Response Function (CADRF) is defined as $\nu(x,t)=E[Y(t)|X=x]$, the expected value over the Individual Dose-Response Functions (IDRF) for all individuals with observed $X$. Similar to Conditional Average Treatment Effects (CATE), to estimate the CADRF we make the following standard assumptions (rubin_causal_2005; hirano_propensity_2004):
  • Unconfoundedness: $Y(t)\perp\!\!\!\perp T|X,\,\forall t\in\mathcal{T}$. This assumption states that, conditional on the observed covariates, the treatment assignment is independent of the potential outcomes. In other words, there are no unobserved confounders that influence both the treatment assignment and the outcome.
  • Overlap or positivity: $0<p(T=t|X=x)<+\infty,\,\forall t\in\mathcal{T}$ with $x\in\mathcal{X}$. The overlap assumption ensures that for every covariate value $x$, there is a positive probability of receiving any treatment level. This is crucial for estimating treatment effects across the entire range of treatment levels.
  • Consistency: $Y=Y(T)$ with probability $1$ with T the observed treatment assignment. This assumption links the observed outcomes to the potential outcomes, stating that the observed outcome is equal to the potential outcome corresponding to the treatment received.
Quantifying the IDRF requires observing the $Y(t)$ for all possible treatment values. These treatment values are all counterfactuals and thus impossible to observe as we only can observe $Y$ for a single treatment value $t$ at a time. Furthermore for estimating the CADRF, likewise with CATE estimation, the distribution of the treatment assignment can bias the estimation (hirano_propensity_2004). This distribution of the treatment assignment is called the propensity distribution, which was initially defined for binary treatments. hirano_propensity_2004 introduced the generalized propensity score (GPS) for continuous treatments that aims to unbias the CATE estimation for continuous treatments. The GPS is defined as $\pi(t_{i}|x)=f_{T|X}(T=t_{i}|X=x)$, which is the evaluation of $T=t_{i}$ on the conditional probability density function $T|X$ (hirano_propensity_2004). The treatment assignment is considered uniformly assigned between lower $t_{L}$ and upper $t_{U}$ possible treatment if $f_{T|X}$ represents the density function of the uniform distribution between $t_{L}$ and $t_{U}$. The GPS can then be used to mimic the randomly assigned treatment to estimate the unbiased CADRF (wu_matching_2024).
@@
The estimated $\widehat{IDRF}$ can also be seen as follows: $\widehat{IDRF}=\nu(x,t)+\epsilon_{a,IDRF}(x,t)+\epsilon_{e,IDRF}(x,t)$. The aleatoric uncertainty, i.e., the data or irreducible uncertainty, is symbolized by $\epsilon_{a,IDRF}(x,t)$ created by the inherent variability between individuals having the same covariates Hullermeier2021_db. $\epsilon_{e,IDRF}(x,t)$ symbolises the epistemic uncertainty, i.e. the reducible uncertainty, coming from model specification and finite samples Hullermeier2021_db. Estimating both uncertainties creates the opportunity to estimate the ranges of the $\widehat{IDRF}$:
Problem Definition: To accurately estimate the $\widehat{IDRF}$ for all possible treatment values we require correctly estimating both uncertainties for all treatment values equally, or more formally; for a specific significance level $\alpha$, lower treatment bound $t_{L}$, upper treatment bound $t_{U}$, and covariates $X$, we require prediction intervals $C(t,X)$ such that
@@ 3 Related Work
Propensity score methods, introduced by rosenbaum_central_1983, have become widespread in causal inference, especially in observational studies. These methods aim to balance confounders across treatment groups, reducing bias in treatment effect estimates. hirano_propensity_2004 generalized this propensity score to continuous instead of binary treatments, introducing the generalized propensity score and building the foundation for causal inference with continuous exposures. wu_matching_2024 used the generalized propensity score for matching continuous treatments to debias the treatment assignment and more accurately estimate the average dose-response curve for all treatment values. Other approaches adapt machine learning techniques to dose-response modelling. For instance, athey_generalized_2019 developed generalized random forests for heterogeneous treatment effect estimation, adaptable to continuous treatments.
@@
To provide UQ, this work adapts conformal prediction. Conformal prediction is a model-agnostic method introduced by vovk_algorithmic_2022 that constructs prediction intervals with guaranteed finite-sample coverage under distribution-free assumptions. Conformal prediction uses conformity scores to assess uncertainty. Various improvements, such as the adaptive version by romano_conformalized_2019, have increased the flexibility and applicability to even heteroscedastic settings. Additionally, lei_distribution_free_2018 and papadopoulos_inductive_2002 introduced split conformal prediction, significantly improving computational efficiency. For scenarios involving covariate or distribution shifts, tibshirani_conformal_2019 introduced weighted conformal prediction to ensure coverage under mismatched training and testing data distributions, with additional work by gibbs_adaptive_2021; gibbs_conformal_nodate and barber_conformal_2023. By reweighting the calibration samples similar to weighted conformal prediction,  guan_localized_2023 introduced localized conformal prediction where the prediction intervals are determined by calibration samples localized around the test sample. vovk_nonparametric_2019 also introduced conformal predictive systems (CPS); an extension of full conformal prediction that allows extracting predictive distributions instead of prediction intervals. More recently, jonkers_conformal_pred_shift_2024 combined previous concepts, introducing weighted conformal predictive systems to also account for covariate shifts.
@@
In causal inference, conformal prediction has mainly been applied to binary treatments. For instance, lei_conformal_2021 were among the first to apply conformal prediction to treatment effects estimation in randomized experiments and confounded or observational data. jonkers_conformal_monte_carlo_2024 and alaa_conformal_nodate extended this approach to the potential outcomes framework, providing uncertainty to quantify individual treatment effects. However, the use of conformal prediction in continuous treatment settings remains largely unexplored. schroder_conformal_2024 proposed a conformal prediction framework for prediction intervals of treatment effects for continuous treatment interventions. However, their approach mainly covers single-treatment interventions and is computationally intensive, requiring optimization per confidence level, treatment, and sample where they provide prediction intervals for a single treatment value. For a more in-depth analysis of schroder_conformal_2024, see Appendix C.
@@ 4.1 Introduction to Conformal Prediction
Before delving into our proposed method, we provide a formal introduction to conformal prediction (jonkers_conformal_pred_shift_2024; tibshirani_conformal_2019). Conformal prediction offers a powerful method for constructing prediction intervals with guaranteed finite-sample coverage under distribution-free assumptions (vovk_algorithmic_2022). The key insight of conformal prediction lies in its use of a nonconformity measure to quantify the degree to which a new observation differs from previously observed data.
@@
for a pre-specified significance level $\alpha\in(0,1)$ where the probability is calculated over the points $Z_{i},i=1,...,n$ .
@@
Finally, we construct the prediction interval containing all $y$ where (jonkers_conformal_pred_shift_2024)
@@
tibshirani_conformal_2019 presented conformal prediction slightly differently by using quantile functions instead, which will be more convenient for weighted conformal prediction later on. tibshirani_conformal_2019 defines the $1-\alpha$ quantile function as follows, where $F_{R}(y)$ represents the distribution of nonconformity scores $R_{i}^{y}$ consisting of a sum of point masses $\delta_{a}$ with mass at $a$ where $R^{y}\sim F_{R}(y)$ (tibshirani_conformal_2019). $F_{R}(y)$ can then be used to calculate probabilities:
@@
This procedure guarantees that $P(y_{n+1}\in\hat{C}(X_{n+1}))\geq 1-\alpha$ for any exchangeable distribution of the data and any choice of nonconformity measure (tibshirani_conformal_2019).
@@ 4.1.1 Inductive Conformal Prediction
The previously mentioned conformal prediction approach is computationally heavy as it requires fitting $n\cdot\#\{\mathbb{R}\}+1$ estimators $\hat{\mu}$. Inductive or split conformal prediction (ICP), introduced by papadopoulos_inductive_2002, tackles this computation issue by splitting the training sequence $Z_{1:n}=\{Z_{1},...,Z_{n}\}$ into two sets: the proper training set $Z_{1:m}=\{Z_{1},...,Z_{m}\}$ and the calibration set $Z_{m+1:n}=\{Z_{m+1},...,Z_{n}\}$. A single regression model $\hat{\mu}$ is fit on the proper training set while the nonconformity scores (e.g., $R_{i}=|y_{i}-\hat{\mu}(X_{i})|,i=m+1,...,n$) are generated from the calibration set. These scores are sorted in descending order denoted as $R_{1}^{*},...,R_{n-m}^{*}$. Then, for a new sample with features $X_{n+1}$, a point prediction is made $\hat{y}_{n+1}=\hat{\mu}(X_{n+1})$. Finally, given a target coverage of $1-\alpha$, the prediction interval becomes
@@
where $s=\lfloor\alpha(n-m+1)\rfloor$ represents the $1-\alpha$ quantile of the ordered nonconformity set with size nmn-m (jonkers_conformal_pred_shift_2024).
@@ 4.1.2 Weighted Conformal Prediction
Evaluating and requiring coverage guarantees for the dose-response model at all possible treatment values changes the test distribution compared to the training distribution. In the training data, all treatment values are sampled according to their (conditional) training distribution, which can be determined by other variables in the case of confounding. However, every treatment value is possible in testing, and thus, every treatment sample can be sampled. This mimics sampling a new test sample with the treatment value from a uniform distribution, which can be vastly different from the treatment distribution in the training data. However, other interventional distributions can also be possible if prior knowledge is available.
Standard conformal prediction only guarantees coverage if the joint distribution of the new sample $Z_{n+1}$ and $Z_{1:n}$ remains the same under permutations, which is called the exchangeability assumption (vovk_algorithmic_2022; tibshirani_conformal_2019). This issue is called covariate shift; The features $X_{n+1}$ come from a different distribution compared to $X_{1:n}$, while the relation between $X$ and $y$ remains the same. More formally: $X_{i}\sim P_{X},\ i=1,...,n$ and $X_{n+1}\sim\tilde{P}_{X}$ where $\tilde{P}_{X}\neq P_{X}$ while $y_{i}\sim P_{Y|X},\ i=1,...,n$.
Weighted conformal prediction provides a solution to tackle this issue (tibshirani_conformal_2019). However, their main assumption is that the likelihood ratio between the training $P_{X}$ and the test covariate distribution $\tilde{P}_{X}$ is known, defined as
@@
The rationale is that they reweight the distribution of nonconformity scores $F_{R}(y)$ to make the nonconformity scores more exchangeable with the test population by using the following weights in equation 7 (tibshirani_conformal_2019):
@@
Consequently, these weights adjust the distribution of nonconformity scores to give more weight to nonconformity scores that are more likely in the test set and vice versa while in standard conformal prediction, every $R_{i}$ has equal weight. Also, note that the weights $p^{w}(x)$ are normalized, cancelling out any constant terms resulting in $w(x)$ being proportional to $w(x)\propto\frac{d\tilde{P}(x)}{dP(x)}$. An extension to split weighted conformal prediction can be done similarly as in section 4.1.1 (tibshirani_conformal_2019).